OpenStudy (anonymous):

why do we say that -1/1=1/-1 if we take square root on both sides it becomes i/1=1/i _______1 and i=1 but actually left and right hand sides of equation 1 are not equal and i is never equal to -1

OpenStudy (kira_yamato):

....

OpenStudy (zarkon):

you are making the assumption that \[\sqrt{\frac{a}{b}}=\frac{\sqrt{a}}{\sqrt{b}}\] holds for all a,b (b not zero). This is not the case

OpenStudy (anonymous):

There is no theorem says a/b = b/a this is equal just because it is specified as 1.